OpenStudy (anonymous):

Look at the following equations: 8P=12-2Q Q+4P=6 Which statement is true about the lines graphed from the equations? A. They coincide. B. They are parallel. C. They are perpendicular. D. They intersect but are not perpendicular.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

try to drawing the graph

OpenStudy (anonymous):

to make a line we need two points

OpenStudy (anonymous):

2P = 12 - 2Q let Q = 1, then we substitute into the equation to obtain P

OpenStudy (anonymous):

\[Q=1\rightarrow 2P = 12 - 2 \left( 1 \right)\]

OpenStudy (anonymous):

Can you continue?

OpenStudy (anonymous):

2P = 12 -2 2P = 10 P = 10/2 P = 5

OpenStudy (anonymous):

I dont think so, still didnt get it

OpenStudy (anonymous):

so we have a point (P, Q) = (1,5)

OpenStudy (anonymous):

to draw a graph, we have to look for two point in the graph

OpenStudy (anonymous):

ah alright

OpenStudy (anonymous):

we already have one point, meaning we are now looking for another point

OpenStudy (anonymous):

try for different Q values

OpenStudy (anonymous):

Can you try it?

OpenStudy (anonymous):

yeah thanks